• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name prout

    Log entry time 20:43:46 on October 21,1998

    Entry number 3241

    keyword=update

    Two NMR measurements have been made this shift 6:00 27.5 - 37.3
    8:00 29.2 - 39.6
    The cell was moved to the empty position at 7:00 in anticipation of tuning beam
    into the Hall but this is now delayed to at least 9:30 when the Compton people
    will do studies. After this the accelerator will take the beam for test plans. Shift
    workers have been sent home. Run 1896 continues to take cosmic data.